OpenStudy (anonymous):

On a game show, there are 10 keys in a bag, and 3 of the keys start a car. A contestant randomly chooses a key. It does not start the car. She returns the key to the bag, the host mixes up the keys, and she randomly selects another key. This key does not start the car either. What is the probability of this no start, no start outcome? (Points : 1) Option A: 49/100 Option B: 7/15 Option C: 21/50 Option D: 9/100

OpenStudy (welshfella):

this is a sequence of independent events so you multiply the individual probabilities

OpenStudy (welshfella):

what is the probability of choosing a key that will not open a car with one pick?

OpenStudy (anonymous):

5

OpenStudy (welshfella):

3/10 is probability for picking a key which opens a car

OpenStudy (anonymous):

take away the 3 that work plus the 2 that they tried gives you 5 so do I multiply 3/10 by 5/10?

OpenStudy (anonymous):

It gave me 3/20 so clearly not what I have to do xD

OpenStudy (welshfella):

no there is only 2 events - 2 picks out of the bag probability of picking wrong key = 1 - 3/10 = 7/10 this is the same for the first pick and the second

OpenStudy (welshfella):

these are multiplied

OpenStudy (welshfella):

no start, no start = 7/10 * 7 /10 = ?

OpenStudy (anonymous):

49/100 But im still a little confused

OpenStudy (welshfella):

49/100 is correct can you understand why the prob for no start = 7/10? there are 7 keys which wont start the car out of a total of 10

OpenStudy (anonymous):

Oooh yea okay I got it. Sorry /.\ Thank you!!!

OpenStudy (welshfella):

yw